# Long-Duration Station-Keeping Strategy for Cislunar Spacecraft Formations

### v1

@@ 1 Introduction
Deployment of Distributed Space Systems (DSS) in cislunar orbits has been gaining interest following renewed attention in the exploration and economic activity of cislunar space. In anticipation of the deployment of NASA’s Gateway Station in a near-rectilinear halo orbit (NRHO) around the Earth-Moon $L_{2}$ libration point [1], development of DSS technologies, including swarming, formation flight, rendezvous, proximity operations, and docking (RPOD) requires further attention. In particular, long-term spacecraft swarming that performs bounded relative motion has the possibility to expand future space mission concepts. However, due to the chaotic nature of the dynamical system around the Moon, traditional guidance, navigation, and control (GN&C) algorithms that are tailored for the relative motion in perturbed Keplerian dynamics (e.g., Low Earth Orbit (LEO), Geostationary Orbit (GEO)) are not directly applicable to GN&C systems for cislunar DSS. For example, the 9:2 Southern $L_{2}$ NRHO, which has received special interest for cislunar missions due to its increased stability, near-constant line-of-sight to Earth, and advantageous position for lunar and scientific operations, still offers numerous unexplored challenges for cislunar DSS, notably fully unstable relative motion.
@@
In LEO, where the majority of distributed space systems have been deployed, the gravitational attraction of the Earth dominates other perturbative accelerations, such as those induced by gravitational harmonics, solar radiation pressure, third body gravitation, and drag. This favorable property means that trajectories in these regimes are not significantly perturbed, allowing for the applicability of low-fidelity models (i.e., the restricted two-body problem), for which semi-analytical solutions exist, and which offer numerous dynamical insights. Indeed, the existence of analytical relative motion models in near-earth orbits relies on the fact that absolute motion can be semi-analytically described [2]. In contrast, low-fidelity models of spacecraft dynamics in cislunar space, such as the CR3BP, have not thus far produced similar levels of dynamical insight for absolute and relative motion. These models, which are non-linear and non-integrable, have not been found to admit analytical solutions. As a result, no exact and analytically available form for relative motion about periodic solutions has been found in these simplified models. Furthermore, simplified models perform significantly worse in describing true spacecraft motion in cislunar space, as a result of the stronger perturbative effects relative to the gravitational attraction of the primaries. For example, the assumption of circularity of the Earth-Moon orbit is a source of significant model error in the CR3BP, since the Moon’s orbit of the Earth is non-negligibly eccentric and wobbling (e 0.0255$\sim$ 0.0255∼ 0.0255 to 0.07750.07750.0775) [3]. Another large source of model error arises from solar gravitation. Inclusions of these effects, namely in the elliptical-restricted three-body problem (ER3BP) or the bi-circular restricted four-body problem (BCR4BP) [4], are still too oversimplified to describe orbital motion in cislunar space with sufficient levels of accuracy necessary for many applications.
@@
Additionally, many orbits of interest in cislunar space are fully unstable (their state transition matrices contain eigenvalues outside of the unit circle) [5], in contrast to Keplerian orbits, which are only marginally unstable (their state transition matrices contain repeated eigenvalues on the unit circle) [6]. This means that spacecraft will diverge from a reference trajectory at an exponential rate (in Keplerian orbits, trajectories diverge at a linear rate). As a result, any spacecraft in these orbits will need to frequently apply absolute and relative station-keeping maneuvers to maintain its orbit. In contrast, many spacecraft missions in LEO which do not demand high levels of absolute positioning accuracy can orbit the Earth for long-durations without requiring any absolute orbit station-keeping.
@@
As a result of the aforementioned perturbative nature of cislunar space, the design of absolute motion in cislunar orbits typically leverages procedures whereby trajectories are first designed in a low-fidelity model, which is easier to work with, provides dynamical insights, and admits periodic solutions, and then converged in a high-fidelity model, which includes perturbative effects and uses an ephemeris to accurately determine the locations of celestial bodies, to remove dynamical infeasibilities. This procedure is commonly used to generate long-duration baseline trajectories in the vicinity of periodic solutions of the CR3BP [4, 7], and is similar to procedures used for some absolute station-keeping techniques in near-Earth orbit, such as for Tandem-X and Terra-SARx [8, 9]. For many orbits of interest in the cislunar regime, this procedure serves as the basis for absolute orbit station-keeping.
@@
Several solution techniques have been proposed for performing absolute orbit control in cislunar orbits [10, 11]. Absolute orbit control is typically divided into long and short horizons [12], wherein long-horizon uncontrolled trajectories are generated in a high-fidelity ephemeris model (HFEM), using the aforementioned procedure of transitioning trajectories, and short-horizon correction maneuvers are performed to clean up dispersions along the long-duration reference trajectory. One of the most popular approaches, known as x-axis crossing [12], has been operated on the CAPSTONE mission [13] and has been studied for use in Gateway station-keeping [12]. It involves solving for the impulsive maneuver that targets an x-axis crossing of the reference trajectory several revolutions from the current time. Other approaches, which all involve an HFEM reference, involve full-state feedback model predictive control [14], chance-constrained feedback control design with missed-thrust events [15], and local eigenmotion control [16].
@@
Relative motion control about periodic orbits of restricted multi-body problems has also been investigated and is typically described through linear dynamical models. In particular, Ref. 17 derives a relative motion model that utilizes local toroidal coordinates to describe spacecraft relative motion and demonstrates its use for station-keeping under high-fidelity dynamics for up to 10 revolutions in the vicinity of a Sun-Earth $L_{1}$ libration point Halo orbit. Ref. 18 extends these techniques by describing relative motion in a velocity-normal-binormal (VNB) frame and applying relevant optimal control techniques to local toroidal coordinates. The VNB frame is also utilized in Ref. 19, in which approach trajectories are generated in modal coordinates by perturbing stable modes of the eigensystem. Other studies simply leverage cartesian coordinates, such as Ref. 20, which demonstrates continuous-thrust linear quadratic regulation in relative dynamics under relative navigation. Refs. 18 and 21 represent one of the few deployments of relative control of swarms under HFEM dynamics, though a key finding of these works is that the controller dynamics model error induces large dispersions and raises $\Delta V$ expenditure.
@@
Though absolute-relative formation-keeping in Keplerian regimes has been demonstrated, such as in the case of TANDEM-X [9, 8], no works have comprehensively formalized a strategy of formation-keeping in multi-body systems that considers the realistic conditions that present considerable difficulty to the success of formation flight in cislunar space. The HELIOSWARM mission [22] involves 9 agent formation flight in a cislunar 2:1 mean motion resonance orbit. However, this orbit offers favorable stability properties and still leverages two-body relative motion models (i.e., the Yamanaka Ankersen STM [23]), meaning many of the station-keeping strategies leveraged by this mission are not extendible to a broad range of cislunar orbits. To the knowledge of the authors, relative orbit generation about NRHOs, which has been performed in simplified dynamical models like the CR3BP, ER3BP, and BCR4BP [21] has not been demonstrated about non-periodic high-fidelity reference trajectories, which allow for significantly lower $\Delta V$ expenditure over long durations. Additionally, perturbative effects to the relative motion induced by maneuver execution error and missed maneuver events in absolute control has not been considered in relative control contexts. Finally, many studies of spacecraft relative motion in cislunar space do not provide guarantees of passive safety under the possibility of missed maneuvers.
@@
This work presents a novel strategy for cislunar formation flight that produces bounded absolute motion about near-rectilinear halo orbits and bounded relative motion of multi-agent formations with guarantees of passive safety. The strategy is simulated under realistic conditions with high-fidelity dynamics, maneuver execution error, missed maneuver events, and navigation uncertainty. By enforcing passive safety with a control barrier function (CBF), the strategy allows for decoupled and modularized computation of the absolute and relative control while efficiently and robustly enforcing formation safety. The resulting strategy produces low total $\Delta V$ expenditure of less than 1 meter per second per agent per year, comparable to $\Delta V$ expenditures of single spacecraft station-keeping.
@@ 2.2 HFEM Dynamics
High-fidelity dynamical models are used to construct long-duration trajectories and verify resulting control schemes. The equations of motion are expressed in a Moon-fixed normalized rotating frame [4]. This results in the equation of motion
@@
Here, $M$ is a direction cosine matrix constructed from ephemeris positions and velocities of the two primaries that describes the rotation from the inertial frame to the synodic frame. $\ddot{\bm{B}}$ is the acceleration of the Moon in the inertial frame, and $\ddot{\bm{R}}$ is the acceleration of the spacecraft in the inertial frame, which can be modeled with gravitational accelerations from the primaries (e.g. Earth, Moon), additional gravitational perturbations (e.g. Sun, Venus), and possibly solar radiation pressure and gravitational harmonics. Relevant ephemeris data is obtained from SPICE [24].
@@ 2.3 Relative Motion Dynamics
Linear relative dynamics can be produced by taking a first order approximation of the absolute dynamics about a reference trajectory. Let $\bm{\phi}(t)$ be a reference trajectory which is a solution to the set of equations of motion $\dot{\bm{x}}(t)=\bm{f}(\bm{x}(t),t)$. If $\bm{\rho}(t)=[\delta\bm{r}(t)^{\top},\delta\bm{v}(t)^{\top}]^{\top}$ is defined with respect to $\bm{\phi}(t)$ then
$\dot{\bm{\rho}}(t)=\bm{f}(\bm{\phi}(t)+\bm{\rho}(t))-\bm{f}(\bm{\phi}(t)).$ (4)
Taking the first-order variation with respect to $\bm{\phi}(t)$ gives
$\dot{\bm{\rho}}(t)\approx A(t)\bm{\rho}(t)$ (5)
@@
This approach to the construction of the relative motion dynamics is agnostic to the absolute dynamics, meaning this approach is valid for both the CR3BP and HFEM dynamics. However, this relative motion must be numerically integrated, and has no analytically solution, unlike for Keplerian regimes [6].
@@ 2.4 Coordinate Frames
Expressing the relative motion in different coordinate frames can often provide additional insight into underlying dynamical structure. Relative motion in Keplerian regimes is often expressed in LVLH frames. A popular approach for cislunar regimes is to express relative motion in a velocity-normal-binormal (VNB) frame (see Refs. 18, 19), where axes are aligned with the along track velocity and moon angular momentum directions (also sometimes referred to as a velocity-normal-conormal (VNC) frame [22]). Let the reference trajectory be decomposed into its positions and velocities, i.e. $\bm{\phi}(t)=[\bm{r}_{syn}(t)^{\intercal},\bm{v}_{syn}(t)^{\intercal}]^{\intercal}$ and let $\bm{r}_{M,syn}$ be the position of the moon with respect to the barycenter. A transformation from a relative position in the synodic frame to the VNB frame can be performed with
$\begin{gathered}\delta\bm{r}_{syn}(t)=R(t)\delta\bm{r}_{VNB}(t),\quad R(t)=\begin{bmatrix}\hat{\bm{x}}&\hat{\bm{y}}&\hat{\bm{z}}\end{bmatrix},\\ \quad\hat{\bm{x}}\parallel\bm{v}_{syn}(t),\quad\hat{\bm{y}}\parallel\left(\bm{v}_{syn}(t)\times(\bm{r}_{syn}(t)-\bm{r}_{M,syn})\right),\quad\hat{\bm{z}}=\hat{\bm{x}}\times\hat{\bm{y}}.\end{gathered}$ (7)
@@ 3.1 Absolute Trajectory
Performing long-duration formation-keeping relies upon the generation of reference trajectories to be tracked by a spacecraft. For absolute motion, this is typically done by taking a periodic orbit in a lower fidelity model (e.g., CR3BP), and performing numerical continuation to a HFEM [12, 4]. Ephemeris trajectories are generated by stacking the apoapsis node of a CR3BP orbit for the desired number of revolutions and converging the initial guess in the HFEM with a multiple shooting scheme. Fig. 1 shows an example of a 50 revolution ephemeris trajectory generated in the vicinity of a 9:2 NRHO from the CR3BP. For more details on trajectory generation in ephemeris models, see Refs. 12, 25.
@@ 3.2 Relative Orbit Reference Generation
The generation of bounded relative motion between multiple spacecraft in Keplerian orbits simply relies on the satisfaction of the energy matching condition, that is that the orbits of interest have equal semi-major axes [26]. For cislunar orbits, generating bounded relative motion can be more challenging, due to the numerical nature of the trajectories. Since periodic orbits exist in one-parameter families in the CR3BP, neighboring periodic orbits will not produce bounded relative motion in general. However, quasi-periodic tori, which exist about periodic solutions, can produce bounded relative motion so long as its period is equivalent with that of the periodic orbit. Because of this, quasi-periodic tori have served as the foundation for generating bounded relative motion in cislunar space [27, 28, 18]. In close proximity to a periodic orbit, these quasi-periodic orbits exist as a linear combination of the eigenvectors associated with unimodular eigenvalues of the Monodromy matrix. The real and imaginary components of these eigenvalues span an invariant circle, which when propagated for a full revolution, repeats on itself, forming a quasi-periodic relative orbit (QPRO). A QPRO about the 9:2 NRHO in the CR3BP along with a selection of relative trajectories on its surface are displayed in Fig. 2 in the VNB frame.
@@
Unlike the CR3BP, ephemeris trajectories are non-periodic, meaning globally bounded relative motion does not exist in general. However, despite the non-periodicity of ephemeris trajectories, their associated state transition matrices do exhibit central modes that are close to unimodular after integer revolutions, which can be perturbed to produce locally bounded relative motion about the trajectory for a finite duration, similar to the procedure used to generate quasi-periodic tori [29, 30]. If $\bm{\phi}(t)$ represents an ephemeris trajectory $\forall t\in T$, and $\Phi(t,t_{0})$ is the associated state transition matrix of the trajectory, then a point in time $t_{1}\in T$ can be identified for which a nearly central mode exists. The eigenvalue and eigenvector associated with this central mode can be denoted by $e_{r}\pm e_{i}i$ and $\bm{w}_{r}\pm\bm{w}_{i}i$. In dynamics that exhibit periodicity (e.g., CR3BP), this central mode is unimodular for the Monodromy matrix and produces an invariant circle. In ephemeris dynamics, an osculating invariant circle can be constructed in a similar manner with

### v2

@@ 1 Introduction
Deployment of Distributed Space Systems (DSS) in cislunar orbits has been gaining interest following renewed attention in the exploration and economic activity of cislunar space. In anticipation of the deployment of NASA’s Gateway Station in a near-rectilinear halo orbit (NRHO) around the Earth-Moon $L_{2}$ libration point [davisgateway], development of DSS technologies, including swarming, formation flight, rendezvous, proximity operations, and docking (RPOD) requires further attention. In particular, long-term spacecraft swarming that performs bounded relative motion has the possibility to expand future space mission concepts. However, due to the chaotic nature of the dynamical system around the Moon, traditional guidance, navigation, and control (GN&C) algorithms that are tailored for the relative motion in perturbed Keplerian dynamics (e.g., Low Earth Orbit (LEO), Geostationary Orbit (GEO)) are not directly applicable to GN&C systems for cislunar DSS. For example, the 9:2 Southern $L_{2}$ NRHO, which has received special interest for cislunar missions due to its increased stability, near-constant line-of-sight to Earth, and advantageous position for lunar and scientific operations, still offers numerous unexplored challenges for cislunar DSS, notably fully unstable relative motion.
@@
In LEO, where the majority of distributed space systems have been deployed, the gravitational attraction of the Earth dominates other perturbative accelerations, such as those induced by gravitational harmonics, solar radiation pressure, third body gravitation, and drag. This favorable property means that trajectories in these regimes are not significantly perturbed, allowing for the applicability of low-fidelity models (i.e., the restricted two-body problem), for which semi-analytical solutions exist, and which offer numerous dynamical insights. Indeed, the existence of analytical relative motion models in near-earth orbits relies on the fact that absolute motion can be semi-analytically described [koenig2017new]. In contrast, low-fidelity models of spacecraft dynamics in cislunar space, such as the CR3BP, have not thus far produced similar levels of dynamical insight for absolute and relative motion. These models, which are non-linear and non-integrable, have not been found to admit analytical solutions. As a result, no exact and analytically available form for relative motion about periodic solutions has been found in these simplified models. Furthermore, simplified models perform significantly worse in describing true spacecraft motion in cislunar space, as a result of the stronger perturbative effects relative to the gravitational attraction of the primaries. For example, the assumption of circularity of the Earth-Moon orbit is a source of significant model error in the CR3BP, since the Moon’s orbit of the Earth is non-negligibly eccentric and wobbling (e $\sim 0.0255$ to $0.0775$) [nasa_moonfactsheet]. Another large source of model error arises from solar gravitation. Inclusions of these effects, namely in the elliptical-restricted three-body problem (ER3BP) or the bi-circular restricted four-body problem (BCR4BP) [parkassessment], are still too oversimplified to describe orbital motion in cislunar space with sufficient levels of accuracy necessary for many applications.
@@
Additionally, many orbits of interest in cislunar space are fully unstable (their state transition matrices contain eigenvalues outside of the unit circle) [zimovan2020near], in contrast to Keplerian orbits, which are only marginally unstable (their state transition matrices contain repeated eigenvalues on the unit circle) [ALFRIEND201083]. This means that spacecraft will diverge from a reference trajectory at an exponential rate (in Keplerian orbits, trajectories diverge at a linear rate). As a result, any spacecraft in these orbits will need to frequently apply absolute and relative station-keeping maneuvers to maintain its orbit. In contrast, many spacecraft missions in LEO which do not demand high levels of absolute positioning accuracy can orbit the Earth for long-durations without requiring any absolute orbit station-keeping.
@@
As a result of the aforementioned perturbative nature of cislunar space, the design of absolute motion in cislunar orbits typically leverages procedures whereby trajectories are first designed in a low-fidelity model, which is easier to work with, provides dynamical insights, and admits periodic solutions, and then converged in a high-fidelity model, which includes perturbative effects and uses an ephemeris to accurately determine the locations of celestial bodies, to remove dynamical infeasibilities. This procedure is commonly used to generate long-duration baseline trajectories in the vicinity of periodic solutions of the CR3BP [parkassessment, zimovanorbitgeneration], and is similar to procedures used for some absolute station-keeping techniques in near-Earth orbit, such as for Tandem-X and Terra-SARx [montenbruck2008navigation, krieger2007tandemx]. For many orbits of interest in the cislunar regime, this procedure serves as the basis for absolute orbit station-keeping.
@@
Several solution techniques have been proposed for performing absolute orbit control in cislunar orbits [shirobokov2017survey, williams2023comparison]. Absolute orbit control is typically divided into long and short horizons [davis2022orbit], wherein long-horizon uncontrolled trajectories are generated in a high-fidelity ephemeris model (HFEM), using the aforementioned procedure of transitioning trajectories, and short-horizon correction maneuvers are performed to clean up dispersions along the long-duration reference trajectory. One of the most popular approaches, known as x-axis crossing [davis2022orbit], has been operated on the CAPSTONE mission [cheetham2022capstone] and has been studied for use in Gateway station-keeping [davis2022orbit]. It involves solving for the impulsive maneuver that targets an x-axis crossing of the reference trajectory several revolutions from the current time. Other approaches, which all involve an HFEM reference, involve full-state feedback model predictive control [shimane2025output], chance-constrained feedback control design with missed-thrust events [shrivastavamarkov], and local eigenmotion control [elangoeigenmotion].
@@
Relative motion control about periodic orbits of restricted multi-body problems has also been investigated and is typically described through linear dynamical models. In particular, Ref. elliott2022phd derives a relative motion model that utilizes local toroidal coordinates to describe spacecraft relative motion and demonstrates its use for station-keeping under high-fidelity dynamics for up to 10 revolutions in the vicinity of a Sun-Earth $L_{1}$ libration point Halo orbit. Ref. takubo2025passively extends these techniques by describing relative motion in a velocity-normal-binormal (VNB) frame and applying relevant optimal control techniques to local toroidal coordinates. The VNB frame is also utilized in Ref. vela2025modal, in which approach trajectories are generated in modal coordinates by perturbing stable modes of the eigensystem. Other studies simply leverage cartesian coordinates, such as Ref. khoury_relative_2024, which demonstrates continuous-thrust linear quadratic regulation in relative dynamics under relative navigation. Refs. takubo2025passively and takubo2025safeoptimalnspacecraftswarm represent one of the few deployments of relative control of swarms under HFEM dynamics, though a key finding of these works is that the controller dynamics model error induces large dispersions and raises $\Delta V$ expenditure.
@@
Though absolute-relative formation-keeping in Keplerian regimes has been demonstrated, such as in the case of TANDEM-X [krieger2007tandemx, montenbruck2008navigation], no works have comprehensively formalized a strategy of formation-keeping in multi-body systems that considers the realistic conditions that present considerable difficulty to the success of formation flight in cislunar space. The HELIOSWARM mission [helioswarm] involves 9 agent formation flight in a cislunar 2:1 mean motion resonance orbit. However, this orbit offers favorable stability properties and still leverages two-body relative motion models (i.e., the Yamanaka Ankersen STM [yamanaka_new_2002]), meaning many of the station-keeping strategies leveraged by this mission are not extendible to a broad range of cislunar orbits. To the knowledge of the authors, relative orbit generation about NRHOs, which has been performed in simplified dynamical models like the CR3BP, ER3BP, and BCR4BP [takubo2025safeoptimalnspacecraftswarm] has not been demonstrated about non-periodic high-fidelity reference trajectories, which allow for significantly lower $\Delta V$ expenditure over long durations. Additionally, perturbative effects to the relative motion induced by maneuver execution error and missed maneuver events in absolute control has not been considered in relative control contexts. Finally, many studies of spacecraft relative motion in cislunar space do not provide guarantees of passive safety under the possibility of missed maneuvers.
@@
This work presents a novel strategy for cislunar formation flight that produces bounded absolute motion about near-rectilinear halo orbits and bounded relative motion of multi-agent formations with guarantees of passive safety. The strategy is simulated under realistic conditions with high-fidelity dynamics, maneuver execution error, missed maneuver events, and navigation uncertainty. By enforcing passive safety with a control barrier function (CBF), the strategy allows for decoupled and modularized computation of the absolute and relative control while efficiently and robustly enforcing formation safety. The resulting strategy produces low total $\Delta V$ expenditure on the order of 1 meter per second per agent per year, comparable to $\Delta V$ expenditures of single spacecraft station-keeping.
@@ 2.2 HFEM Dynamics
High-fidelity dynamical models are used to construct long-duration trajectories and verify resulting control schemes. The equations of motion are expressed in a Moon-fixed normalized rotating frame [parkassessment]. This results in the equation of motion
@@
Here, $M$ is a direction cosine matrix constructed from ephemeris positions and velocities of the two primaries that describes the rotation from the inertial frame to the synodic frame. $\ddot{\bm{B}}$ is the acceleration of the Moon in the inertial frame, and $\ddot{\bm{R}}$ is the acceleration of the spacecraft in the inertial frame, which can be modeled with gravitational accelerations from the primaries (e.g. Earth, Moon), additional gravitational perturbations (e.g. Sun, Venus), and possibly solar radiation pressure and gravitational harmonics. Relevant ephemeris data is obtained from SPICE [actonspice].
@@ 2.3 Relative Motion Dynamics
Linear relative dynamics can be produced by taking a first order approximation of the absolute dynamics about a reference trajectory. Let $\bm{\phi}(t)$ be a reference trajectory which is a solution to the set of equations of motion $\dot{\bm{x}}(t)=\bm{f}(\bm{x}(t),t)$. If $\delta\bm{x}(t)=[\delta\bm{r}(t)^{\intercal},\delta\bm{v}(t)^{\intercal}]^{\intercal}$ is defined with respect to $\bm{\phi}(t)$ then
$\delta\dot{\bm{x}}(t)=\bm{f}(\bm{\phi}(t)+\delta\bm{x}(t))-\bm{f}(\bm{\phi}(t)).$ (4)
Taking the first-order variation about $\delta\bm{x}(t)=0$ gives
$\delta\dot{\bm{x}}(t)\approx A(t)\delta\bm{x}(t)$ (5)
@@
This approach to the construction of the relative motion dynamics is agnostic to the absolute dynamics, meaning this approach is valid for both the CR3BP and HFEM dynamics. However, this relative motion must be numerically integrated, and has no analytical solution, unlike for Keplerian regimes [ALFRIEND201083].
@@ 2.4 Coordinate Frames
Expressing the relative motion in different coordinate frames can often provide additional insight into underlying dynamical structure. Relative motion in Keplerian regimes is often expressed in LVLH frames. A popular approach for cislunar regimes is to express relative motion in a velocity-normal-binormal (VNB) frame (see Refs. takubo2025passively, vela2025modal), where axes are aligned with the along track velocity and moon angular momentum directions (also sometimes referred to as a velocity-normal-conormal (VNC) frame [helioswarm]). Let the reference trajectory be decomposed into its positions and velocities, i.e. $\bm{\phi}(t)=[\bm{r}^{\intercal}_{syn}(t),\bm{v}^{\intercal}_{syn}(t)]^{\intercal}$ and let $\bm{r}_{M,syn}$ be the position of the moon with respect to the system origin. A transformation from a relative position in the synodic frame to the VNB frame can be performed with
[…]
@@ 3.1 Absolute Trajectory
Performing long-duration formation-keeping relies upon the generation of reference trajectories to be tracked by a spacecraft. For absolute motion, this is typically done by taking a periodic orbit in a lower fidelity model (e.g., CR3BP), and performing numerical continuation to a HFEM [davis2022orbit, parkassessment]. Ephemeris trajectories are generated by stacking the apoapsis node of a CR3BP orbit for the desired number of revolutions and converging the initial guess in the HFEM with a multiple shooting scheme. Fig. 1 shows an example of a 50 revolution ephemeris trajectory generated in the vicinity of a 9:2 NRHO from the CR3BP. For more details on trajectory generation in ephemeris models, see Refs. davis2022orbit, zimovanorbitgeneration.
@@ 3.2 Relative Orbit Reference Generation
The generation of bounded relative motion between multiple spacecraft in Keplerian orbits simply relies on the satisfaction of the energy matching condition, that is, the orbits of interest have equal semi-major axes [schaubspacecraftformationflying]. For cislunar orbits, generating bounded relative motion can be more challenging, due to the numerical nature of the trajectories. Since periodic orbits exist in one-parameter families in the CR3BP, neighboring periodic orbits will not produce bounded relative motion in general. However, quasi-periodic tori, which exist about periodic solutions, can produce bounded relative motion so long as its period is equivalent with that of the periodic orbit. Because of this, quasi-periodic tori have served as the foundation for generating bounded relative motion in cislunar space [elliott2022describing, baresi2017spacecraft, takubo2025passively]. In close proximity to a periodic orbit, these quasi-periodic orbits exist as a linear combination of the eigenvectors associated with unimodular eigenvalues of the Monodromy matrix. The real and imaginary components of these eigenvalues span an invariant circle, which when propagated for a full revolution, repeats on itself, forming a quasi-periodic relative orbit (QPRO). A QPRO about the 9:2 NRHO in the CR3BP along with a selection of relative trajectories on its surface are displayed in Fig. 2 in the VNB frame.
@@
Unlike the CR3BP, ephemeris trajectories are non-periodic, meaning globally bounded relative motion does not exist in general. However, despite the non-periodicity of ephemeris trajectories, their associated state transition matrices do exhibit central modes that are close to unimodular after integer revolutions, which can be perturbed to produce locally bounded relative motion about the trajectory for a finite duration, similar to the procedure used to generate quasi-periodic tori [olikara2016computation, foss2024thesis]. If $\bm{\phi}(t)$ represents an ephemeris trajectory $\forall t\in T$, and $\Phi(t,t_{0})$ is the associated state transition matrix of the trajectory, then a point in time $t_{1}\in T$ can be identified for which a nearly central mode exists. The eigenvalue and eigenvector associated with this central mode can be denoted by $e_{r}\pm e_{i}i$ and $\bm{w}_{r}\pm\bm{w}_{i}i$. In dynamics that exhibit periodicity (e.g., CR3BP), this central mode is unimodular for the Monodromy matrix and produces an invariant circle. In ephemeris dynamics, an osculating invariant circle can be constructed in a similar manner with